# Can LLMs Generate Tabular Summaries of Science Papers? Rethinking the Evaluation Protocol

### v3

@@ Abstract
Literature review tables are essential for summarizing and comparing collections of scientific papers. We explore the task of generating tables that best fulfill a user’s informational needs given a collection of scientific papers. Building on recent work Newman et al. (2024), we extend prior approaches to address real-world complexities through a combination of LLM-based methods and human annotations. Our contributions focus on three key challenges encountered in real-world use: (i) User prompts are often under-specified; (ii) Retrieved candidate papers frequently contain irrelevant content; and (iii) Task evaluation should move beyond shallow text similarity techniques and instead assess the utility of inferred tables for information-seeking tasks (e.g., comparing papers). To support reproducible evaluation, we introduce arXiv2Table, a more realistic and challenging benchmark for this task, along with a novel approach to improve literature review table generation in real-world scenarios. Our extensive experiments on this benchmark show that both open-weight and proprietary LLMs struggle with the task, highlighting its difficulty and the need for further advancements. Our dataset and code are available at https://github.com/JHU-CLSP/arXiv2Table.
Can LLMs Generate Tabular Summaries of Science Papers?
Rethinking the Evaluation Protocol
Weiqi Wang♣♠, Jiefu Ou, Yangqiu Song, Benjamin Van Durme, Daniel Khashabi Center for Speech and Language Processing, Johns Hopkins University, Baltimore, MD, USA Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China {wwang194, jou6, vandurme, danielk}@jhu.edu, {wwangbw, yqsong}@cse.ust.hk
@@ 1 Introduction
With recent advancements in large language models (LLMs; OpenAI, 2025b; DeepSeek-AI et al., 2025), several studies Newman et al. (2024); Dagdelen et al. (2024); Sun et al. (2024) have explored generating literature review tables by prompting LLMs with a set of pre-selected papers and the table’s caption. While these efforts represent meaningful progress, we argue that the existing task definition and evaluation protocols are somewhat unrealistic, thus hindering the practical applicability of generation methods.
@@
In this paper, we introduce our task, as illustrated in Figure 1, which improves upon previous task definitions through two key adaptations. First, our pilot study shows that LLMs struggle to retrieve relevant papers from large corpora. To benchmark this, we introduce distractor papers by selecting them based on semantic similarity to papers in the ground-truth table. LLMs must first determine which papers should be included before generating the table. Second, we replace table captions with abstract user demands that describe the goal of curating the table, making the task more aligned with real-world scenarios. We build upon the ArxivDigesTables Newman et al. (2024) dataset and construct a sibling benchmark through human annotation to verify the selected distractors, comprising 1,957 tables and 7,158 papers.
Meanwhile, current evaluation methods rely on static semantic embeddings to estimate schema overlap between generated and ground-truth tables and require human annotations to assess the quality of unseen schemas and values. However, semantic embeddings struggle to capture nuanced, context-specific variations due to their reliance on pre-trained representations, while human annotation is costly and time-consuming. Moreover, the most effective table generation approaches define schemas primarily based on paper abstracts. This method risks missing important aspects present in the full text, leading to loosely defined schemas with inconsistent granularity.
@@ 2 Related Works
Prior works primarily attempt to generate scientific tables through two stages: schema induction and value extraction. For schema induction, early methods like entity-based table generation Zhang and Balog (2018) focused on structured input, while recent work has explored schema induction from user queries Wang et al. (2024) and comparative aspect extraction Hashimoto et al. (2017). For value extraction, various approaches such as document-grounded question-answering Kwiatkowski et al. (2019); Dasigi et al. (2021); Lee et al. (2023), aspect-based summarization Ahuja et al. (2022), and document summarization DeYoung et al. (2021); Lu et al. (2020) have been proposed to extract relevant information. Beyond these methods, several datasets have been introduced to support scientific table-related tasks, such as TableBank Li et al. (2020), SciGen Moosavi et al. (2021), and SciTabQA Lu et al. (2023). Recently, Newman et al. (2024) proposed streamlining schema and value generation with LLMs sequentially and curated a large-scale benchmark for evaluation. However, all these methods assume a clean and fully relevant set of papers and rely on predefined captions or abstract-based schemas, which risk missing key details. In contrast, we argue for an evaluation approach where candidate papers include tangentially relevant or distracting papers, aligning more closely with real-world literature review workflows Padmakumar et al. (2025).
@@ 3 Task Definition
We first define a pipeline consisting of three subtasks that extend prior definitions and better capture the real-world usage of literature review tabular generation. For all the following tasks, we are given a user demand prompt $p$, which specifies the intended purpose of creating the table. (T1) Candidate Paper Retrieval: We begin with a given universe of papers (e.g., the content of Google Scholar or arXiv) from which relevant papers need to be identified. Given a large collection, the goal is to use a search engine (IR) to retrieve a subset of candidate papers $C:=\{d_{i}\}_{i=1}^{M}$ of size $M$, which may include distractor papers—i.e., papers that resemble the user demand prompt but do not fully satisfy the requirement. (T2) Paper Selection: Given $C$, the second subtask is to select the relevant subset of size $m$ ($m<M$): $R:=\{d_{i}\}_{i=1}^{m}\subseteq C$, which best aligns with the user demand $p$. T2 differs from T1 in scale. Due to the large scale of T1, IR engines must optimize for recall, ensuring that as many relevant papers as possible are retrieved. However, T2 operates at a smaller scale, where precision is the priority, as it focuses on filtering out distractors and selecting only the most relevant papers. (T3) Table Induction: Given the selected papers $R$, the objective is to generate a table with $m$ rows and $N$ columns, where $N\geq 2$ (i.e., no single-column tables). Each row $r_{i}\in\{r_{1},r_{2},\dots,r_{m}\}$ corresponds to a unique input document $d_{i}\in R$, and each column $c_{j}\in\{c_{1},c_{2},\dots,c_{N}\}$ represents a unique aspect of the documents. We refer to these $N$ columns as the schema of the table and the $N\times m$ cells as the values of the table. The value of each cell is derived from its respective document according to the aspect defined by the corresponding column.
@@ 4.1 Constructing User Demand Prompts
The first step is to collect user demands $p$ that explicitly describe the desired table (can be understood without the table content) and do not reveal the table’s schema or specific values.
@@
While the input dataset contains one caption per table, collected from arXiv papers, these captions are meant to complement tables rather than fully describe them. As a result, they are generally concise. For example, a table caption might read: “Performance comparison of different approaches,” which is too vague to understand without seeing the table. Consequently, using table captions as prompts may not yield a well-defined task. A more contextually self-contained rewritten user demand might instead be: “Draft a table that compares different knowledge editing methods, focusing on their performance on QA datasets.”
@@
To address this issue, we propose rewriting the captions of literature review tables into abstract yet descriptive user intentions using LLMs. We guide GPT-4o with a prompt (see §A) that first explains the task to the LLM, specifying that the user demand should be sufficiently contextualized to clearly state the table’s purpose while avoiding the inclusion or direct description of column names or specific values. GPT-4o is then expected to infer the user demand for the given table and its caption. Here, LLM is used solely for rewriting existing table captions into user demand prompts and for generating QA pairs grounded in ground-truth tables. These reformulations are strictly tied to observed data and do not require external factual knowledge, minimizing risks of contamination or model-specific bias. For simplicity, we collect only one user demand per table. More examples are provided in Appendix C.
Table captions vs. constructed user demand prompts
To verify that our collected user demands align with our objective, we visualize: (1) the distribution of the number of tokens in the original and modified user demands, and (2) the ratio of captions and user demands of different lengths that have token overlap with the schema or values. From Figure 2, we observe that our modified user demands are generally longer than the original captions, providing a more detailed description of the table’s goal. Furthermore, as shown in Table 1, user demands exhibit a significantly lower overlap ratio with the schema and table values, resulting in fewer overlapping tokens.
4.2 Paper Retrieval Simulation
The unreliability of paper retrieval
We observe consistently low precision and recall across different retrieval sizes, highlighting the challenge of retrieving relevant papers from a noisy corpus. This demonstrates that the first subtask is non-trivial and may introduce noise into subtask T2. However, various information retrieval engines, such as Google Scholar and Semantic Scholar, can replace LMs in this subtask. Thus, we decide to simulate T1 by manually adding noisy distractor papers into $C$ to construct $R$, ensuring a noisy input for T2. This allows us to focus on evaluating LLMs’ capabilities in the T2 and T3 subtasks.
Similarity-based paper retrieval
Moving forward, we associate distractor paper candidates with each table to simulate a potentially noisy document pool before constructing the table. Ideally, distractor candidates should be semantically related to the table but exhibit key differences that fail to meet the user demand. To select such candidates, we adopt a retrieve-then-annotate approach. First, we use a SentenceBERT encoder $F$ to obtain embeddings for (1) the user demand $F(p)$ and (2) all papers in the corpus $\{F(d_{i})\mid d_{i}\in C\}$. Each paper’s embedding is computed by encoding the concatenation of its title and abstract. We then rank all papers $d_{i}\notin R$ based on the average of two cosine similarities: (1) the similarity between the candidate and the user demand, and (2) the average similarity between the candidate and each referenced paper:
@@ 4.2 Paper Retrieval Simulation
After selecting these candidates, we conduct human annotations to verify whether they should indeed be excluded from the table. Given that annotating these tables requires expert knowledge in computer science, we recruit seven postgraduate students with research experience in the field as annotators. To ensure they are well-prepared for the task, the annotators undergo rigorous training, including pilot annotation exams. Their task is to make a binary decision on whether a given distractor paper—based on its title, abstract, user demand, the ground-truth table, and the titles and abstracts of all referenced papers—should be included in the table. Each table contains annotations for 10 papers, with each distractor paper initially assigned to two randomly selected annotators. If both annotators agree on the label, it is finalized. Otherwise, two additional annotators review the paper until a consensus is reached. In the first round, the inter-annotator agreement (IAA) is 94% based on pairwise agreement, and the Fleiss’ Kappa Fleiss (1971) score is 0.73, indicating a substantial level of agreement Landis and Koch (1977). Finally, for each table, we randomly select a number of distractor papers between $[m,10]$.
@@ 4.3 Evaluation via LLM-based Utilization
We guide GPT-4o in generating binary QA pairs based on the ground-truth table. For the first two aspects, we generate QA pairs for all columns and cells, whereas for the third, we randomly sample 10 cell pairs per table and synthesize them into QA pairs. We then prompt GPT-4o to answer these questions based on the generated table, providing yes/no responses. If the answer cannot be found, the model is instructed to respond with no,” and vice versa for yes.” The ratio of “yes” answers indicates how well the generated table preserves the schema, individual values, and pairwise relationships. This represents the recall of the ground-truth table, measuring how much original information is retained in the generated table.
@@
To additionally evaluate precision, we reverse the process: instead of generating QA pairs from the ground-truth table, we generate them from the generated table and ask another LLM to answer them using the ground-truth table. The precision score reflects how much of the generated table’s content is actually supported by the original data. By computing the ratio of “yes” answers, we quantify the accuracy of the generated table in reflecting genuine ground-truth information, as well as any additional useful information not present in the ground-truth table.
@@ 5.1 Baseline Methods
To overcome both issues, Newman et al. Newman et al. (2024) introduces a two-stage process. In the first stage, the model selects papers relevant to the user demand based on their titles and abstracts, then generates a corresponding schema. In the second stage, the model loops through the selected papers and fills in the respective rows based on the full text of each document. A minor drawback of this method is that the schema is generated solely from titles and abstracts, which may overlook details present only in the full text. Note that this method is the strongest recent baseline for scientific tabular generation while other text-to-table methods Deng et al. (2024b) are not directly applicable due to different assumptions.

### v5

# arXiv2Table: Toward Realistic Benchmarking and Evaluation for LLM-Based Literature-Review Table Generation

@@ Abstract
[…]
arXiv2Table: Toward Realistic Benchmarking and Evaluation for LLM-Based Literature-Review Table Generation
Weiqi Wang♣♠, Jiefu Ou, Yangqiu Song, Benjamin Van Durme, Daniel Khashabi Center for Speech and Language Processing, Johns Hopkins University, Baltimore, MD, USA Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China {wwangbw, yqsong}@cse.ust.hk, {jou6, vandurme, danielk}@jhu.edu
@@ 1 Introduction
With recent advancements in large language models (LLMs; OpenAI, 2025c; DeepSeek-AI et al., 2025), several studies Newman et al. (2024); Dagdelen et al. (2024); Sun et al. (2024) have explored generating literature review tables by prompting LLMs with a set of pre-selected papers and the table’s caption. While these efforts represent meaningful progress, we argue that the existing task definition and evaluation protocols are somewhat unrealistic, thus hindering the practical applicability of generation methods.
@@
In this paper, we introduce our task, as illustrated in Figure 1, which improves upon previous task definitions through two key adaptations. First, our pilot study shows that LLMs struggle to retrieve relevant papers from large corpora. To benchmark this, we introduce distractor papers by selecting them based on semantic similarity to papers in the ground-truth table. LLMs must first determine which papers should be included before generating the table. Second, we simulate well-specified yet schema-agnostic user demands that describe the goal of curating the table without revealing column labels or cell values. This formulation is more realistic than terse captions while avoiding schema/value leakage that would bias evaluation. We build upon the ArxivDigesTables Newman et al. (2024) dataset and construct a sibling benchmark through human annotation to verify the selected distractors, comprising 1,957 tables and 7,158 papers. Meanwhile, current evaluation methods rely on static semantic embeddings to estimate schema overlap between generated and ground-truth tables and require human annotations to assess the quality of unseen schemas and values. However, semantic embeddings struggle to capture nuanced, context-specific variations due to their reliance on pre-trained representations, while human annotation is costly and time-consuming. Moreover, the most effective table generation approaches define schemas primarily based on paper abstracts. This method risks missing important aspects present in the full text, leading to loosely defined schemas with inconsistent granularity.
@@ 2 Related Works
Prior works primarily attempt to generate scientific tables through two stages: schema induction and value extraction. For schema induction, early methods like entity-based table generation Zhang and Balog (2018) focused on structured input, while recent work has explored schema induction from user queries Wang et al. (2024b) and comparative aspect extraction Hashimoto et al. (2017). For value extraction, various approaches such as document-grounded question-answering Kwiatkowski et al. (2019); Dasigi et al. (2021); Lee et al. (2023), aspect-based summarization Ahuja et al. (2022), and document summarization DeYoung et al. (2021); Lu et al. (2020) have been proposed to extract relevant information.
Beyond these methods, several datasets have been introduced to support scientific table-related tasks, such as TableBank Li et al. (2020), SciGen Moosavi et al. (2021), and SciTabQA Lu et al. (2023). Ramu et al. (2024) propose an entailment-oriented evaluation complementary to our QA-based protocol. Recently, Newman et al. (2024) proposed streamlining schema and value generation with LLMs sequentially and curated a large-scale benchmark for evaluation. However, all these methods assume a clean and fully relevant set of papers and rely on predefined captions or abstract-based schemas In contrast, we argue for an evaluation approach where candidate papers include tangentially relevant or distracting papers, aligning more closely with real-world literature review workflows Padmakumar et al. (2025).
@@ 3 Task Definition
We first define a pipeline consisting of three subtasks that extend prior definitions and better capture the real-world usage of literature review tabular generation. For all the following tasks, we are given a user demand prompt $p$, which specifies the intended purpose of creating the table.
(T1) Candidate Paper Retrieval: We begin with a given universe of papers (e.g., the content of Google Scholar or arXiv) from which relevant papers need to be identified. Given a large collection, the goal is to use a search engine (IR) to retrieve a subset of candidate papers $C:=\{d_{i}\}_{i=1}^{M}$ of size $M$, which may include distractor papers—i.e., papers that resemble the user demand prompt but do not fully satisfy the requirement.
(T2) Paper Selection: Given $C$, the second subtask is to select the relevant subset of size $m$ ($m<M$): $R:=\{d_{i}\}_{i=1}^{m}\subseteq C$, which best aligns with the user demand pp. T2 differs from T1 in scale. Due to the large scale of T1, IR engines must optimize for recall, ensuring that as many relevant papers as possible are retrieved. However, T2 operates at a smaller scale, where precision is the priority, as it focuses on filtering out distractors and selecting only the most relevant papers.
(T3) Table Induction: Given the selected papers $R$, the objective is to generate a table with $m$ rows and NN columns, where $N\geq 2$ (i.e., no single-column tables). Each row $r_{i}\in\{r_{1},r_{2},\dots,r_{m}\}$ corresponds to a unique input document $d_{i}\in R$, and each column $c_{j}\in\{c_{1},c_{2},\dots,c_{N}\}$ represents a unique aspect of the documents. We refer to these NN columns as the schema of the table and the $N\times m$ cells as the values of the table. The value of each cell is derived from its respective document according to the aspect defined by the corresponding column.
@@ 4.1 Constructing User Demand Prompts
We simulate well-specified yet schema-agnostic user demands $p$ by rewriting original survey-table captions into user-demand-style prompts that better reflect how a researcher would request a comparison table. Each rewritten prompt is required to be (i) self-contained, meaning it is understandable without seeing the table, (ii) goal-oriented, meaning it clearly states the purpose of the table, and (iii) schema/value non-leaking, meaning it does not include gold column names, specific cell values, or direct paraphrases of them.
@@
While the input dataset contains one caption per table, collected from arXiv papers, these captions are meant to complement tables rather than fully describe them. As a result, they are generally concise. For example, a table caption might read: “Performance comparison of different approaches,” which is too vague to understand without seeing the table. Consequently, using table captions as prompts may not yield a well-defined task. A more contextually self-contained rewritten user demand might instead be: “Draft a table that compares different knowledge editing methods, focusing on their performance on QA datasets.” Operationally, a caption is deemed “under-specified” when it cannot unambiguously determine a comparison goal without seeing either the gold schema or table body.
@@
To address this issue, we rewrite the captions of literature review tables into abstract yet descriptive user intentions using GPT-4o. We guide the model with a prompt (see §B) that explains the rewriting task and specifies that the resulting user demand should be sufficiently contextualized to clearly state the table’s purpose while avoiding the inclusion or direct description of column names or specific values. Here, the LLM is used solely for rewriting existing table captions into user demand prompts and for generating QA pairs grounded in ground-truth tables. These reformulations are strictly tied to observed data and do not require external factual knowledge, minimizing risks of contamination or model-specific bias.
To enforce our constraints, we apply an automatic leak check immediately after rewriting (Appendix B) and discard or rewrite any demand that reveals schema or value information. We also perform manual spot checks during construction to remove under-specified requests or prompts that implicitly reveal gold schema/value tokens. For simplicity, we collect only one user demand per table. More examples are provided in Appendix D.
@@
Table captions vs. constructed user demand prompts
To verify that our collected user demands align with our objective, we visualize: (1) the distribution of the number of tokens in the original and modified user demands, and (2) the ratio of captions and user demands of different lengths that have token overlap with the schema or values. From Figure 2, we observe that our modified user demands are generally longer than the original captions, providing a more detailed description of the table’s goal. Furthermore, as shown in Table 1, user demands exhibit a significantly lower overlap ratio with the schema and table values, resulting in fewer overlapping tokens. Quantitative statistics on lexical leakage and overlap are reported in the main paper.
4.2 Paper Retrieval Simulation
The unreliability of paper retrieval
We observe consistently low precision and recall across different retrieval sizes, highlighting the challenge of retrieving relevant papers from a noisy corpus. This demonstrates that the first subtask is non-trivial and may introduce noise into subtask T2. However, various information retrieval engines, such as Google Scholar and Semantic Scholar, can replace LMs in this subtask. Thus, we decide to simulate T1 by adding human-verified distractor papers into the candidate pool $C$, yielding a noisy input for T2 (paper selection on $C$ to produce $R$). This allows us to focus on evaluating LLMs’ capabilities in the T2 and T3 subtasks.
Similarity-based paper retrieval
Moving forward, we associate distractor paper candidates with each table to simulate a potentially noisy document pool before constructing the table. Ideally, distractor candidates should be semantically related to the table but exhibit key differences that fail to meet the user demand. To select such candidates, we adopt a retrieve-then-annotate approach. First, we use a SentenceBERT encoder $F$ to obtain embeddings for (1) the user demand $F(p)$ and (2) all papers in the corpus $\{F(d_{i})\mid d_{i}\in U\}$. Each paper’s embedding is computed by encoding the concatenation of its title and abstract. We then rank all papers $d_{i}\notin R$ based on the average of two cosine similarities: (1) the similarity between the candidate and the user demand, and (2) the average similarity between the candidate and each referenced paper:
@@ 4.2 Paper Retrieval Simulation
After selecting these candidates, we conduct human annotations to verify whether they should indeed be excluded from the table. Given that annotating these tables requires expert knowledge in computer science, we recruit a trained team of annotators with research experience in the field as annotators. To ensure they are well-prepared for the task, the annotators undergo rigorous training, including pilot annotation exams. Their task is to make a binary decision on whether a given distractor paper—based on its title, abstract, user demand, the ground-truth table, and the titles and abstracts of all referenced papers—should be included in the table.
Each table contains annotations for 10 papers, with each distractor paper initially assigned to two randomly selected annotators. If both annotators agree on the label, it is finalized. Otherwise, two additional annotators review the paper until a consensus is reached. In the first round, the inter-annotator agreement (IAA) is 94% based on pairwise agreement, and the Fleiss’ Kappa Fleiss (1971) score is 0.73, indicating a substantial level of agreement. Finally, for each table, we randomly select a number of distractor papers between $[m,10]$ and merge them with $R$ to form $C$.
@@ 4.3 Evaluation via LLM-based Utilization
We guide GPT-4o in generating binary QA pairs based on the ground-truth table. For the first two aspects, we generate QA pairs for all columns and cells, whereas for the third, we randomly sample 10 cell pairs per table and synthesize them into QA pairs. We then prompt GPT-4o to answer these questions based on the generated table, providing yes/no responses. If the answer cannot be found, the model is instructed to respond with “no”, and vice versa for “yes”. The ratio of “yes” answers indicates how well the generated table preserves the schema, individual values, and pairwise relationships. This represents the recall of the ground-truth table, measuring how much original information is retained in generations.
@@
To additionally evaluate precision, we reverse the process: instead of generating QA pairs from the ground-truth table, we generate them from the generated table and ask another LLM (by default, we use the same backbone for consistency; cross-model swaps are reported in Appendix A.3) to answer them using the ground-truth table. The precision score reflects how much of the generated table’s content is actually supported by the original data. By computing the ratio of “yes” answers, precision quantifies how much of the generated table is supported by the ground-truth table; novel content beyond gold is not credited under this automatic protocol. To assess evaluator dependence, we swap QA synthesizer/answerer model families and observe stable method rankings; results are in Appendix A.3.
@@ 5.1 Baseline Methods
To overcome both issues, Newman et al. (2024) introduce a two-stage process. In the first stage, the model selects papers relevant to the user demand based on their titles and abstracts, then generates a corresponding schema. In the second stage, the model loops through the selected papers and fills in the respective rows based on the full text of each document. A minor drawback of this method is that the schema is generated solely from titles and abstracts, which may overlook details present only in the full text. Note that this method is the strongest recent baseline for scientific tabular generation while other text-to-table methods Deng et al. (2024b) are not directly applicable due to different assumptions.
To probe whether explicit reasoning mitigates multi-step synthesis errors, we also evaluate COT-augmented variants that add lightweight chain-of-thought style scaffolding to the strongest baseline and to our method.
Backbone Model Method Paper (T2) Schema Unary Value Pairwise Value Avg
R P F1 P R F1 P R F1 P R F1
LLaMa-3.3 (70B) Baseline 1 52.8 50.0 51.4 31.3 37.7 34.2 29.6 40.4 34.2 28.4 31.8 30.0 32.8
Baseline 2 65.4 63.0 64.2 26.7 69.3 38.5 17.0 56.8 26.2 11.2 22.5 15.0 26.6
Newman et al. 61.9 60.0 60.9 36.4 40.5 38.3 32.8 44.5 37.8 29.5 30.2 29.8 35.3
Ours 69.3 66.5 67.9 41.9 55.4 47.7 43.1 62.6 51.1 36.4 46.9 41.0 46.6
Mistral-Large (123B) Baseline 1 54.7 51.5 53.0 33.1 34.5 33.8 31.6 30.4 31.0 15.5 24.7 19.0 27.9
Baseline 2 66.8 64.0 65.4 27.4 65.0 38.5 22.7 47.4 30.7 17.8 30.7 22.6 30.6
Newman et al. 67.9 65.5 66.7 39.9 41.6 40.7 34.7 46.3 39.7 29.9 35.1 32.3 37.6
Ours 71.3 68.0 69.6 45.4 56.7 50.4 43.3 61.5 50.8 42.0 49.2 45.3 48.8
DeepSeek-V3 (685B) Baseline 1 57.5 55.0 56.2 38.7 41.7 40.1 32.5 43.8 37.3 28.7 31.8 30.1 35.8
Baseline 2 69.8 67.0 68.4 34.9 69.0 46.4 27.1 55.5 36.4 25.7 32.7 28.8 37.2
Newman et al. 70.9 68.5 69.7 39.4 44.2 41.7 36.6 49.2 42.0 33.3 36.5 34.8 39.5
Ours 74.3 71.0 72.6 39.6 56.9 46.7 47.7 65.2 55.1 40.4 49.8 44.6 48.8
GPT-4o-mini Baseline 1 55.9 53.0 54.4 32.0 35.7 33.7 28.9 39.3 33.3 25.0 31.0 27.7 31.6
Baseline 2 68.2 65.0 66.6 31.5 67.7 43.0 27.7 50.8 35.9 21.6 28.3 24.5 34.5
Newman et al. 69.3 66.0 67.6 40.3 45.9 42.9 38.3 47.5 42.4 35.0 37.8 36.3 40.5
Ours 72.6 69.0 70.8 46.5 59.7 52.3 49.0 66.7 56.5 43.5 51.9 47.3 52.0
GPT-4o Baseline 1 58.5 56.0 57.2 35.8 43.2 39.2 36.9 41.8 39.2 29.0 34.7 31.6 36.7
Baseline 2 70.2 67.0 68.6 34.2 68.0 45.5 27.9 56.0 37.2 19.4 33.6 24.6 35.8
Newman et al. 71.3 68.5 69.9 45.0 47.9 46.4 38.7 49.8 43.6 36.9 40.0 38.4 42.8
Newman et al. + COT 72.5 47.0 49.0 48.0 40.0 51.0 45.0 39.0 42.0 40.5 44.5
Ours 74.6 71.5 73.0 51.5 59.4 55.2 46.1 66.7 54.5 45.9 55.7 50.3 53.3
Ours + COT 75.8 53.0 60.5 56.5 48.0 68.0 56.3 47.0 57.0 51.5 55.0
GPT-o3 Ours 77.2 55.0 62.0 58.3 50.0 70.0 58.3 49.0 59.0 53.5 56.7
Table 2: Tabular evaluation results (%) on arXiv2Table. All tasks use P/R/F1 as evaluation metrics. The best within each backbone is underlined and the best overall is bold.